OpenStudy (anonymous):

simplify: square root of 50*x^3*y^4

OpenStudy (anonymous):

\[\sqrt{50x^3y^4} =\sqrt{2.5.5.x.x.x.y.y.y.y}\\ =\sqrt{2.5^2.x.x^2.y^2.y^2}\\=5.x.y.y \sqrt{2.x}\\ =5xy^2\sqrt{2x}\ Ans.\]

OpenStudy (anonymous):

thank you!